# Tree Search for Language Model Agents

### v1

@@ Abstract
Autonomous agents powered by language models (LMs) have demonstrated promise in their ability to perform decision-making tasks such as web automation. However, a key limitation remains: LMs, primarily optimized for natural language understanding and generation, struggle with multi-step reasoning, planning, and using environmental feedback when attempting to solve realistic computer tasks. Towards addressing this, we propose an inference-time search algorithm for LM agents to explicitly perform exploration and multi-step planning in interactive web environments. Our approach is a form of best-first tree search that operates within the actual environment space, and is complementary with most existing state-of-the-art agents. It is the first tree search algorithm for LM agents that shows effectiveness on realistic web tasks. On the challenging VisualWebArena benchmark, applying our search algorithm on top of a GPT-4o agent yields a 39.7% relative increase in success rate compared to the same baseline without search, setting a state-of-the-art success rate of 26.4%. On WebArena, search also yields a 28.0% relative improvement over a baseline agent, setting a competitive success rate of 19.2%. Our experiments highlight the effectiveness of search for web agents, and we demonstrate that performance scales with increased test-time compute. We conduct a thorough analysis of our results to highlight improvements from search, limitations, and promising directions for future work. Our code and models are publicly released at jykoh.com/search-agents.
@@ 1 Introduction
Our experiments show that this search procedure is complementary with existing LM agents, and enables these models to perform better on harder and longer horizon tasks. On VisualWebArena (Koh et al., 2024), search improves the performance of a baseline GPT-4o (OpenAI, 2024) agent by 39.7% relative to the baseline without search, setting a new state-of-the-art (SOTA) success rate of 26.4%. On WebArena (Zhou et al., 2024b), search is also highly effective, contributing a 28.0% relative improvement (yielding a competitive success rate of 19.2%). We also demonstrate that search benefits from scale: allowing the agent to leverage increased test-time compute to improve performance. Our code and models are publicly released at jykoh.com/search-agents.
@@ 2.1 Realistic Simulated Web Environments
Towards the goal of developing autonomous web agents powered by large language models, several prior works focused on building evaluation benchmarks for measuring the progress of models on web tasks. Mind2Web (Deng et al., 2023) is an evaluation benchmark that measures the ability of frontier models in predicting actions taken on static Internet pages. VisualWebBench (Liu et al., 2024) introduced a multimodal benchmark for assessing the ability of models to understand web content. Others have looked towards simulators (as opposed to static HTML content): MiniWoB (Shi et al., 2017; Liu et al., 2018) was one of the first interactive simulators for web tasks, but consisted of simplified environments that do not directly translate into real world performance. WebShop (Yao et al., 2022a) simulates a simplified e-commerce site with real world data. WebLINX (Lù et al., 2024) proposes a benchmark for tackling conversational web navigation, which involves communication between the agent and a human instructor. MMInA (Zhang et al., 2024b) and OSWorld (Xie et al., 2024a) propose benchmarks to measure the ability of agents to accomplish tasks by navigating across multiple computer applications and webpages. WorkArena (Drouin et al., 2024) is a benchmark and simulated environment for tasks on the ServiceNow platform. WebArena (WA) (Zhou et al., 2024b) is a benchmark of 812 tasks across 5 realistic self-hosted re-implementations of popular websites (Shopping, Reddit, CMS, GitLab, Maps), each populated with real world data. VisualWebArena (VWA) (Koh et al., 2024) is a multimodal extension to WebArena, consisting of 910 new tasks across realistic re-implementations of 3 popular real world sites (Classifieds, Reddit, Shopping). To solve tasks in VWA, agents must leverage visual grounding and understand image inputs, providing a realistic and challenging test for multimodal agents.
@@
As the (V)WA environments are one of the most realistic and comprehensive evaluation suites for web tasks, we primarily benchmark our method on (V)WA. We briefly describe the setting here but refer readers to Zhou et al. (2024b) for additional context. The environment $\mathcal{E}=(\mathcal{S},\mathcal{A},\Omega,T)$ consists of a set of states $S$, actions $A$ (Tab. 1), and a deterministic transition function $T:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ that defines transitions between states conditioned on actions. Each task in the benchmark consists of a goal specified with a natural language instruction $I$ (e.g., “Find me the cheapest red Toyota car below \$2000.”). Each task has a predefined reward function $R:\mathcal{S}\times\mathcal{A}\rightarrow\{0,1\}$ which measures whether an agent’s execution is successful. We implement our search algorithm on the (V)WA web simulators, but our method is fully general and can be applied to any setting with an interactive environment that the agent can freely explore.
@@ 2.2 Language-Guided Autonomous Agents
Autonomous web agents, powered by frontier (multimodal) language models (Google, 2023; OpenAI, 2024; Anthropic, 2024), are the SOTA approaches for many of the above benchmarks. Kim et al. (2024) showed that large language models can be prompted to execute computer tasks on MiniWoB++ (Liu et al., 2018), requiring far fewer demonstrations than reinforcement learning methods. AutoWebGLM (Lai et al., 2024) collects web browsing data for curriculum training and develops a web navigation agent based off a 6B parameter language model that outperforms GPT-4 on WebArena. Patel et al. (2024) showed that a language model agent can improve its performance through finetuning on its own synthetically generated data. Pan et al. (2024) show that introducing an automatic evaluator to provide guidance on task failure or success can improve the performance of a baseline Reflexion (Shinn et al., 2024) agent. Fu et al. (2024) extracts domain knowledge from offline data and provides this to the language agent during inference, to enable it to leverage helpful domain knowledge. Sodhi et al. (2024) proposes a method to enable agents to dynamically compose policies to solve a diverse set of web tasks. Our procedure is an inference-time approach that is compatible with many of these past approaches that develop better base agents.
In the multimodal setting, WebGUM (Furuta et al., 2024) finetuned a 3B parameter multimodal language model on a large corpus of demonstrations, achieving strong performance on MiniWoB and WebShop. Koh et al. (2024) showed that prompting multimodal language models with a Set-of-Marks (Yang et al., 2023a) representation enables the model to navigate complex webpages more effectively than text-only agents. SeeAct (Zheng et al., 2024) demonstrated that frontier multimodal models such as GPT-4V (Yang et al., 2023b) and Gemini (Google, 2023) can be grounded and prompted to follow natural language instructions for automating web tasks.
@@ 2.3 Search and Planning
In contrast to prior work, our setting is grounded in realistic web environments, and we search over the actual environment space (i.e., the web). This means that the search mechanics need to incorporate not just the text outputs of the agent, but also external environmental feedback from a highly complex (and realistic) environment. Our experiments show that our environmentally grounded tree search substantially improves the performance of language model agents.
@@ 3.1 Agent Backbone
Most SOTA web agents are built through prompting large (multimodal) language models (Zhou et al., 2024b; Pan et al., 2024; Fu et al., 2024; Zheng et al., 2024; Koh et al., 2024). A pretrained language model or multimodal model $f_{\phi}$ is prompted with the current webpage observation $o_{t}$ and instructed to predict the next action $a_{t}$ to be executed. It is common to leverage prompting techniques, such as ReAct (Yao et al., 2022b), RCI (Kim et al., 2024), or Chain-of-Thought (CoT) prompting (Wei et al., 2022), to improve the performance of the agent. Language model agents also allow us to sample a diverse set of actions (e.g., with nucleus sampling (Holtzman et al., 2020)), which is essential for creating plausible branches to explore during search (see Sec. 3.3). Our proposed search algorithm can in principle be applied to any baseline language agent model. We show in Sec. 4 that search improves inference-time performance on a range of models without retraining or finetuning $f_{\phi}$.
@@ 3.2 Value Function
We implement a best-first search heuristic using a value function $f_{v}$ which estimates the expected reward $\mathbb{E}[R(s_{t})]$ of the current state $s_{t}$, where the ground truth goal state would provide a perfect reward of 1. As the state $s_{t}$ of the simulator is not always accessible to the agent ($s_{t}$ may include private information such as database entries of the site), the value function computes the value $v_{t}$ using the current and previous observations of the agent, as well as the natural language task instruction $I$:
@@ 3.3 Search Algorithm
Our proposed search algorithm is a best-first search method loosely inspired by A* search (Hart et al., 1968), a classic graph traversal algorithm used widely in computer science. We use a language model agent to propose candidate branches of the search tree. The search has hyperparameters depth $d$, branching factor $b$, and search budget $c$ which determine the maximum size of the search tree, [^1] In Sec. 5.1 we show that increasing the size of the search tree improves results at the expense of using increased compute for exploration. and termination threshold $\theta$. We describe the procedure in detail in the following paragraphs and provide the formal algorithm in Appendix A.3.
@@
At each iteration of the search process, we extract the next item from the frontier, $s_{p}\leftarrow\text{pop}(\mathcal{F})$. We use the value function to compute the score for state $s_{p}$ (with observation $o_{p}$ and previous observations $o_{1},\ldots,o_{p-1}$):
@@
If $v_{p}\geq\theta$ (the agent is likely to have found a goal state) or $s\geq c$ (the search budget has been exceeded), we will terminate the search and navigate to the best state $\hat{s}_{t}$ found thus far.
Otherwise, if the current branch does not exceed the maximum depth (i.e., $|(s_{0},\ldots,s_{p})|<d$), we will generate plausible next actions for branching by obtaining $b$ candidate actions $\{a_{p}^{1},\ldots,a_{p}^{b}\}$ from the language model agent $f_{\phi}$. For each $i$, we execute $a_{p}^{i}$ and add the resulting state $s_{p}^{i}$ to the frontier with the score of the current state [^2] We opt for this approach instead of immediately computing the value for resulting states $s_{p}^{i}$ as immediate evaluation requires more backtracking calls, which would incur much more overhead in the (V)WA simulators. This is because the deterministic environments allow the partially observable states to be represented by the sequences of actions that reached them.:
@@ 4.1 Implementation Details
Our search algorithm is compatible with most off-the-shelf language model agents. In this work, we test it with simpler, more general, prompt-based agents, and leave incorporation of our method with more performant methods that incorporate domain-specific techniques (Fu et al., 2024; Sodhi et al., 2024) for future work. We run several agent baselines (full prompts provided in the appendix):
  • GPT-4o + SoM: We run the multimodal GPT-4o (OpenAI, 2024) (gpt-4o-2024-05-13) based agent with the same prompt from Koh et al. (2024). We similarly apply a preprocessing step to assign a Set-of-Marks (SoM) (Yang et al., 2023a) representation to the webpage. This highlights every interactable element on the webpage with a bounding box and a unique ID. The input to the agent is a screenshot of the SoM-annotated webpage, and a text description of the elements on the page with their corresponding SoM IDs.
  • Llama-3-70B-Instruct: We run the caption-augmented Llama-3-70B-Instruct agent with the same prompt from Koh et al. (2024). We generate captions for each image on the webpage using an off-the-shelf captioning model (in our case, BLIP-2; Li et al. 2023). The accessibility tree [^3] https://developer.mozilla.org/en-US/docs/Glossary/Accessibility_tree representation of the webpage observation is extracted and provided to the language model as its input observation at each step.
  • GPT-4o: On WebArena (which does not require visual grounding), we run a text-only GPT-4o agent using the same prompt from Zhou et al. (2024b). Similar to the Llama-3 baseline, this model uses an accessibility tree representation of the current webpage as its input observation (but this baseline does not include image captions).
@@
As detailed in Sec. 3.2, we require a value function which scores the likelihood that the current state $s_{t}$ is a goal state. We implement the value function by prompting a multimodal language model (specifically, GPT-4o (OpenAI, 2024) (gpt-4o-2024-05-13)) with the task instruction $I$, screenshots of the agent’s trajectory, previous actions the agent took, and the current page URL. The full prompt is provided in Appendix A.2.2. The multimodal LM is instructed to output whether the current state is a success, a failure, and if it’s a failure, whether it is on a trajectory towards success. These outputs are assigned values of 1, 0, and 0.5 respectively (and 0 for invalid output). In order to get more finegrained and reliable scores, we leverage ideas from self-consistency prompting (Wang et al., 2023), and sample multiple reasoning paths by prompting the multimodal LM with CoT (Wei et al., 2022). We sample 20 different paths from the GPT-4o model using ancestral sampling (temperature of 1.0 and top-$p$ of 1.0). The final value assigned to state $s_{t}$, used in the best-first search heuristic, is computed by averaging the values from each of the 20 reasoning paths. In our implementation, sampling from the value function is significantly less costly than predicting the next action, as action prediction requires more input tokens for few-shot examples and the text representation of the page. We estimate the total API cost of the GPT-4o agent for predicting the next action to be approximately 2×2$\times$2 × that of computing the value of a state.
@@ 4.2 Results
Benchmark Agent Model Max Actions No Search + Search $\Delta$
VisualWebArena Llama-3-70B-Instruct (Koh et al., 2024) 30 9.8% - -
@@
Llama-3-70B-Instruct (ours) 5 7.6% 16.7% +119.7%
GPT-4o + SoM (ours) 18.9% 26.4% +39.7%
WebArena GPT-4o (Zhou et al., 2024b) 30 13.1% - -
GPT-4o (Zhou et al., 2024b) 13.1% - -
@@
GPT-4o (ours) 5 15.0% 19.2% +28.0%
Table 2: Success rates (SR) and relative change ($\Delta$) for baseline models and models that employ search on the VisualWebArena (Koh et al., 2024) and WebArena (Zhou et al., 2024b) benchmarks. We also show other published approaches. Search substantially improves our baseline models, setting a new state-of-the-art on VisualWebArena.
@@
With the Llama-3 caption-augmented agent on VWA, we see that search brings even more significant improvements, more than doubling the success rate. With search, success rate improves by 119.7% relative to the baseline (7.6% to 16.7%). We attribute this to the GPT-4o model used for the value function being a generally stronger (multimodal) model than Llama-3. With search, Llama-3-70B-Instruct achieves success rates that are close to the best frontier multimodal models. As Llama-3 has openly available model weights, the strong performance of the Llama-3-70B-Instruct agent with search can prove to be a cost effective agent model for iteration and development in future work which requires access to model internals.

### v2

@@ Abstract
Autonomous agents powered by language models (LMs) have demonstrated promise in their ability to perform decision-making tasks such as web automation. However, a key limitation remains: LMs, primarily optimized for natural language understanding and generation, struggle with multi-step reasoning, planning, and using environmental feedback when attempting to solve realistic computer tasks. Towards addressing this, we propose an inference-time search algorithm for LM agents to explicitly perform exploration and multi-step planning in interactive web environments. Our approach is a form of best-first tree search that operates within the actual environment space, and is complementary with most existing state-of-the-art agents. It is the first tree search algorithm for LM agents that shows effectiveness on realistic web tasks. On the challenging VisualWebArena benchmark, applying our search algorithm on top of a GPT-4o agent yields a 39.7% relative increase in success rate compared to the same baseline without search, setting a state-of-the-art success rate of 26.4%. On WebArena, search also yields a 28.0% relative improvement over a baseline agent, setting a competitive success rate of 19.2%. Our experiments showcase the effectiveness of search for web agents, and we demonstrate that performance scales with increased test-time compute. We conduct a thorough analysis of our results to highlight improvements from search, limitations, and promising directions for future work. Our code and models are publicly released at jykoh.com/search-agents.
@@ 1 Introduction
Our experiments show that this search procedure is complementary with existing LM agents, and enables these models to perform better on harder and longer horizon tasks. On VisualWebArena (Koh et al., 2024), search improves the performance of a baseline GPT-4o (OpenAI, 2024) agent by 39.7% relative to the baseline without search, setting a new state-of-the-art (SOTA) success rate of 26.4%. On WebArena (Zhou et al., 2024b), search is also highly effective, contributing a 28.0% relative improvement (yielding a competitive success rate of 19.2%). We also demonstrate that our search procedure benefits from scale: achieving improved performance as the agent is alloted greater amounts of test-time computation. Our code and models are publicly released at jykoh.com/search-agents.
@@ 2.1 Realistic Simulated Web Environments
Towards the goal of developing autonomous web agents powered by large language models, several prior works focused on building evaluation benchmarks for measuring the progress of models on web tasks. Mind2Web (Deng et al., 2023) is an evaluation benchmark that measures the ability of frontier models in predicting actions taken on static Internet pages. VisualWebBench (Liu et al., 2024b) introduced a multimodal benchmark for assessing the ability of models to understand web content. Others have looked towards simulators (as opposed to static HTML content): MiniWoB (Shi et al., 2017; Liu et al., 2018) was one of the first interactive simulators for web tasks, but consisted of simplified environments that do not directly translate into real world performance. WebShop (Yao et al., 2022a) simulates a simplified e-commerce site with real world data. WebLINX (Lù et al., 2024) proposes a benchmark for tackling conversational web navigation, which involves communication between the agent and a human instructor. MMInA (Zhang et al., 2024b) and OSWorld (Xie et al., 2024a) propose benchmarks to measure the ability of agents to accomplish tasks by navigating across multiple computer applications. WorkArena (Drouin et al., 2024) is a simulated environment for tasks on the ServiceNow platform. WebArena (WA) (Zhou et al., 2024b) is a benchmark of 812 tasks across 5 realistic self-hosted re-implementations of popular websites (Shopping, Reddit, CMS, GitLab, Maps), each populated with real world data. VisualWebArena (VWA) (Koh et al., 2024) is a multimodal extension to WebArena, consisting of 910 new tasks across realistic re-implementations of 3 popular real world sites (Classifieds, Reddit, Shopping). To solve tasks in VWA, agents must leverage visual grounding and understand image inputs, providing a realistic and challenging test for multimodal agents.
@@
As the (V)WA environments are one of the most realistic and comprehensive evaluation suites for web tasks, we primarily benchmark our method on (V)WA. We briefly describe the setting here but refer readers to Zhou et al. (2024b) for additional context. The environment $\mathcal{E}=(\mathcal{S},\mathcal{A},\Omega,T)$ consists of a set of states $S$, actions $A$ (Tab. 1), and a deterministic transition function $T:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ that defines transitions between states conditioned on actions. Each task in the benchmark consists of a goal specified with a natural language instruction $I$ (e.g., “Find me the cheapest red Toyota car below \$2000.”). Each task has a predefined reward function $R:\mathcal{S}\times\mathcal{A}\rightarrow\{0,1\}$ which measures whether an agent’s execution is successful. We implement our search algorithm on the (V)WA web simulators, but our method is fully general and can be applied to any setting with an interactive environment.
@@ 2.2 Language-Guided Autonomous Agents
Autonomous web agents, powered by frontier (multimodal) language models (Google, 2023; OpenAI, 2024; Anthropic, 2024), are the SOTA approaches for many of the above benchmarks. Kim et al. (2024) showed that large language models can be prompted to execute computer tasks on MiniWoB++ (Liu et al., 2018), requiring far fewer demonstrations than reinforcement learning methods. AutoWebGLM (Lai et al., 2024) collects web browsing data for curriculum training and develops a web navigation agent based off a 6B parameter language model that outperforms GPT-4 on WebArena. Patel et al. (2024) showed that a language model agent can improve its performance through finetuning on its own synthetically generated data. Pan et al. (2024) show that introducing an automatic evaluator to provide guidance on task failure or success can improve the performance of a baseline Reflexion (Shinn et al., 2024) agent. Fu et al. (2024) extracts domain knowledge from offline data and provides this to the language agent during inference, to enable it to leverage helpful domain knowledge. SteP (Sodhi et al., 2024) and AWM (Wang et al., 2024) propose methods to enable agents to dynamically compose policies to solve web tasks.
In the multimodal setting, WebGUM (Furuta et al., 2024) finetuned a 3B parameter multimodal language model on a large corpus of demonstrations, achieving strong performance on MiniWoB and WebShop. Koh et al. (2024) showed that prompting multimodal language models with a Set-of-Marks (Yang et al., 2023a) representation enables the model to navigate complex webpages more effectively than text-only agents. SeeAct (Zheng et al., 2024) demonstrated that frontier multimodal models such as GPT-4V (Yang et al., 2023b) and Gemini (Google, 2023) can be grounded and prompted to solve web tasks. ICAL (Sarch et al., 2024) builds a memory of multimodal insights from demonstrations and human feedback, improving performance on VisualWebArena.
Our procedure is an inference-time approach that is compatible with many of these past approaches that focus on developing better base agents.
@@ 2.3 Search and Planning
In contrast to prior work, our setting is grounded in realistic web environments, and we search over the actual environment space (i.e., the web). This means that the search mechanics need to incorporate not just the text outputs of the agent, but also external environmental feedback from a highly complex environment.
@@ 3.1 Agent Backbone
Most SOTA web agents are built through prompting large (multimodal) language models (Zhou et al., 2024b; Pan et al., 2024; Fu et al., 2024; Zheng et al., 2024; Koh et al., 2024). A pretrained language model or multimodal model $f_{\phi}$ is prompted with the current webpage observation $o_{t}$ and instructed to predict the next action $a_{t}$ to be executed. It is common to leverage prompting techniques, such as ReAct (Yao et al., 2022b), RCI (Kim et al., 2024), or Chain-of-Thought (CoT) prompting (Wei et al., 2022), to improve the performance of the agent. Language model agents also allow us to sample a diverse set of actions (e.g., with nucleus sampling (Holtzman et al., 2020)), which is essential for creating plausible branches to explore during search (see Sec. 3.3). Our proposed search algorithm can in principle be applied to any base agent. We show in Sec. 4 that search improves inference-time performance on a range of models without retraining or finetuning $f_{\phi}$.
@@ 3.2 Value Function
We implement a best-first search heuristic using a value function $f_{v}$ which estimates the expected reward $\mathbb{E}[R(s_{t})]$ of the current state $s_{t}$, where the ground truth goal state would provide perfect reward of 1. As the state $s_{t}$ of the simulator is not always accessible to the agent ($s_{t}$ may include private information such as database entries of the site), the value function computes the value $v_{t}$ using the current and previous observations, as well as the natural language task instruction $I$:
@@ 3.3 Search Algorithm
Our proposed search algorithm is a best-first search method loosely inspired by A* search (Hart et al., 1968), a classic graph traversal algorithm used widely in computer science. We use a language model agent to propose candidate branches of the search tree. The search has hyperparameters depth $d$, branching factor $b$, and search budget $c$ which determine the maximum size of the search tree, [^1] In Sec. 5.1 we show that increasing the size of the search tree improves results at the expense of using increased compute for exploration. and termination threshold $\theta$. The search procedure is summarized in Fig. 1. We describe it in detail in the following paragraphs and provide the formal algorithm in Appendix A.4.
@@
At each iteration of the search process, we extract the next state from the frontier, $s_{p}\leftarrow\text{pop}(\mathcal{F})$. We use the value function to compute the score for state $s_{p}$ (with observation $o_{p}$ and previous observations $o_{1},\ldots,o_{p-1}$):
@@
If $v_{p}\geq\theta$ (i.e., the agent is likely to have found a goal state) or $s\geq c$ (the search budget has been exceeded), we will terminate the search and navigate to the best state $\hat{s}_{t}$ found thus far. Otherwise, if the current branch does not exceed the maximum depth (i.e., $|(s_{0},\ldots,s_{p})|<d$), we will generate plausible next actions for branching by obtaining $b$ candidate actions $\{a_{p}^{1},\ldots,a_{p}^{b}\}$ from the language model agent $f_{\phi}$. For each $i$, we execute $a_{p}^{i}$ and add the resulting state $s_{p}^{i}$ to the frontier with the score of the current state [^2] We opt for this approach instead of immediately computing the value for resulting states $s_{p}^{i}$ as immediate evaluation requires more backtracking calls, which would incur much more overhead in the (V)WA simulators.:
@@ 4.1 Implementation Details
Our search algorithm is compatible with most off-the-shelf language model agents. In this work, we test it with simpler, more general, prompt-based agents, and leave incorporation of our method with more performant methods that incorporate domain-specific techniques (Fu et al., 2024; Sodhi et al., 2024) for future work. We run several prompt-based agent baselines with different input formats (full prompts provided in the appendix):
  • Multimodal SoM: For multimodal models that accept multiple image-text inputs, such as GPT-4o (OpenAI, 2024) (gpt-4o-2024-05-13), we run the multimodal agent from Koh et al. (2024) with the same prompt. We similarly apply a preprocessing step to assign a Set-of-Marks (SoM) (Yang et al., 2023a) representation to the webpage. This highlights every interactable element on the webpage with a bounding box and a unique ID. The input to the agent is a screenshot of the SoM-annotated webpage, and a text description of the elements on the page with their corresponding SoM IDs.
  • Caption-augmented: For base models that are not multimodal (e.g., Llama-3-70B-Instruct (Dubey et al., 2024)), we run the caption-augmented agent with the same prompt from Koh et al. (2024). We generate captions for each image on the webpage using an off-the-shelf captioning model (in our case, BLIP-2; Li et al. 2023). The accessibility tree [^3] https://developer.mozilla.org/en-US/docs/Glossary/Accessibility_tree representation of the webpage observation is extracted and provided to the language model as its input observation at each step.
  • Text-only: On WebArena (which does not require visual grounding), we run text-only agents using the prompt from Zhou et al. (2024b), for both GPT-4o and Llama-3-70B-Instruct. Similar to the caption-augmented baseline, this model uses an accessibility tree representation of the current webpage as its input observation (but this text-only baseline does not include image captions).
@@
As detailed in Sec. 3.2, we require a value function which scores the likelihood that the current state $s_{t}$ is a goal state. We implement the value function by prompting a multimodal language model with the task instruction $I$, screenshots of the agent’s trajectory, previous actions the agent took, and the current page URL. The full prompt is provided in Appendix A.3.2. The multimodal LM is instructed to output whether the current state is a success, a failure, and if it’s a failure, whether it is on a trajectory towards success. These outputs are assigned values of 1, 0, and 0.5 respectively (and 0 for invalid output). In order to get more finegrained and reliable scores, we leverage ideas from self-consistency prompting (Wang et al., 2023), and sample multiple reasoning paths by prompting the multimodal LM with CoT (Wei et al., 2022). We sample 20 different paths from the GPT-4o model using ancestral sampling (temperature of 1.0 and top-$p$ of 1.0). The final value assigned to state $s_{t}$, used in the best-first search heuristic, is computed by averaging the values from each of the 20 reasoning paths. In our implementation, calling the value function is significantly cheaper than predicting the next action, as action prediction consumes more input tokens for few-shot examples and the representation of the page. We estimate the API cost of the GPT-4o SoM agent for action prediction to be approximately 2×2$\times$2 × that of computing the value.
@@ 4.2 Results
Agent Model Max Steps No Search + Search $\Delta$
VWA Llama-3-70B-Instruct + captions (Koh et al., 2024) 30 9.8% - -
@@
ICAL (Sarch et al., 2024) 23.4% - -
Llama-3-70B-Instruct + captions 5 7.6% 16.7% +119.7%
GPT-4o + SoM 18.9% 26.4% +39.7%
WA GPT-4o (Zhou et al., 2024b) 30 13.1% - -
@@
Llama-3-70B-Instruct 5 7.6% 10.1% +32.3%
GPT-4o 15.0% 19.2% +28.0%
Table 2: Success rates (SR) and relative change ($\Delta$) for baseline models and models that employ search on the VisualWebArena (VWA) (Koh et al., 2024) and WebArena (WA) (Zhou et al., 2024b) benchmarks. We also show other published approaches. Search substantially improves our baseline models, setting a new state-of-the-art on VWA.
@@
With weaker base models, we also observe substantial improvements. For the Llama-3 caption-augmented agent on VWA, introducing search improves the success rate on VWA by 119.7% relative to the baseline (7.6% to 16.7%). With search, Llama-3-70B-Instruct achieves success rates that are close to the best frontier multimodal models that do not use search. On WebArena, we also see a substantial relative improvement of 32.2% for the text-based Llama-3 agent (7.6% to 10.1%). The strong performance of the Llama-3-70B-Instruct agent with search can prove to be a cost effective agent model for iteration in future work that requires access to model internals. These results over a variety of model scales and capabilities demonstrate the generality and effectiveness of our approach.